OpenStudy (anonymous):

The equation of line CD is y = -2x - 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5).

OpenStudy (anonymous):

y = -2x + 13 y = -1/2x + 7 y = -1/2x + 3 y = − 2x − 3

OpenStudy (anonymous):

I know that the slope has to be the same.

OpenStudy (anonymous):

So either A or D

OpenStudy (anonymous):

parallel lines have the same slope... then sub in pt. (4,5) now you know that m=-2, x=4, and y=5...sub. into y=mx+b to find b

OpenStudy (anonymous):

to i have to input..

OpenStudy (anonymous):

substitute m=-2, x=4, and y=5 to y=mx+b to find b

OpenStudy (anonymous):

im going with D y = − 2x − 3 is this correct..

OpenStudy (anonymous):

did you solve it?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

what did you get as b?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

how would i do it?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

substitute m=-2, x=4, and y=5 to y=mx+b to find b

OpenStudy (anonymous):

b = 13

OpenStudy (anonymous):

@AntiNode

OpenStudy (anonymous):

so its A?

OpenStudy (anonymous):

@FlvsGirl

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks a lot i have 2 more easy questions if you can help

OpenStudy (anonymous):

did you understand how to get it?

OpenStudy (anonymous):

ok I will try

OpenStudy (anonymous):

The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (-3, 2). y = 5x + 13 y = -1/5 x + 7/5 y = - 1/5 x -7/5 y = 5x - 17

OpenStudy (anonymous):

what will be the slope here?

OpenStudy (anonymous):

-1/5

OpenStudy (anonymous):

now you have m=-1/5, x=-3, y=2... as what we did before substitute these values in y=mx+b

OpenStudy (anonymous):

b = 7/5

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

y = -1/5 x + 7/5

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have one more @AntiNode :-)

OpenStudy (anonymous):

ok..I may reply late because OS is lagging